OpenStudy (baldymcgee6):

If b > 0, how can we write a number a as a logarithm in base b?

OpenStudy (anonymous):

\(a = \log_b b^a\)

OpenStudy (anonymous):

if thats what u looking for

OpenStudy (baldymcgee6):

okay, but why?

OpenStudy (anonymous):

you familiar with inverse functions ?

OpenStudy (baldymcgee6):

yes

OpenStudy (anonymous):

exponent and log are inverses of each other

OpenStudy (anonymous):

u send 'a' to exponent u take log of above result u get ur 'a' back

OpenStudy (anonymous):

\(f(x) = b^x\) \(g(x) = \log x\) if g is the inverse of f, g(f(x)) = x

OpenStudy (anonymous):

g(f(a)) = a

OpenStudy (anonymous):

g(b^a) = a

OpenStudy (baldymcgee6):

okay, thank you

OpenStudy (anonymous):

log_b b^a = a wlcme :)